$SPAD/src/input XPOLY.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

)cl all
poly := XPolynomial(Integer)
pr: poly := 2*x + 3*y-5
pr2: poly := pr*pr
pd  := expand pr
pd2 := pd*pd
expand(pr2) - pd2
qr :=  pr**3
qd :=  pd**3
 trunc(qd,2)
trunc(qr,2)
Word := OrderedFreeMonoid Symbol
w: Word := x*y**2
rquo(qr,w)
sh(pr,w::poly)

———-

References^{[1] nothing}

4